# Towards Robust Incremental Learning under Ambiguous Supervision

### v3

@@ Introduction
To reduce annotation costs, we study a novel weakly supervised learning framework dubbed incremental partial label learning (IPLL), where (i)-data is given sequentially as a stream; (ii)-each task contains a vast number of new class samples while potentially carries old class data, and (iii)-each sample is associated with a coarse candidate label set instead of the ground truth; see Figure 1. Notably, since the annotator may confuse previous experience with the true label of the sample, our IPLL setup allows the candidate label set of the sample to include both new and old classes, while ensuring the inclusion of the true one. Arguably, the IPLL problem is deemed more practical in real-world scenarios thanks to its relatively lower cost to annotations.
The key to successful learning from ambiguous supervision is label disambiguation, i.e., identifying the true labels from the candidate sets. To achieve this, existing partial label learning (PLL) algorithms (Wang et al. 2022c; Yan and Guo 2023; Li et al. 2023; Jia et al. 2024) mostly rely on the self-training algorithm that assigns pseudo-labels using the model predictions. However, such a strategy can be problematic in our IPLL framework. On one hand, the dynamically changing environment amplifies knowledge confusion that hinders the disambiguation procedure. The forgetting of past knowledge intensifies the bias of deep neural networks (DNNs). On the other hand, ambiguous supervision significantly undermines the ability of DNNs to acquire meaningful knowledge from samples, exacerbating catastrophic forgetting. As exemplified in our empirical studies, directly combining the PLL method with the IL method also leads to inferior results. Specifically, when combining the most popular PLL method PiCO (Wang et al. 2022c) with one classic IL method iCaRL (Rebuffi et al. 2017), the model overly emphasizes old classes in the early stages of each task (see Figure 2(a)), thus exacerbating the confusion of old class knowledge (see Figure 2(b)). In the more challenging Tiny-ImageNet, the negative impacts of classification bias become more pronounced (see Figure 6(c)). This gives rise to the fundamental issue in IPLL—how to balance the model’s perception of new and old knowledge.
@@
Figure 2: (a) The real/estimated distribution of old classes (0$\sim$9) and new classes (10$\sim$19) in the early stage of the second task of the CIFAR100 (10 tasks). Combining the PLL baseline (PiCO) and the IL baseline (iCaRL) leads to classification bias. (b) Comparison of our PGDR and PiCO with iCaRL in IPLL. It displays the average accuracy of new and old classes for each task in CIFAR100 where PiCO with iCaRL demonstrates inferior and unstable performance compared to our method.
We propose the Prototype-Guided Disambiguation and Replay Algorithm (PGDR), which leverages class prototypes carrying rich category information as proxies to balance the model’s perception of classes. Concretely, to alleviate label confusion, we propose a prototype-guided label disambiguation strategy, which firstly performs distance-based old/new sample separation and then, momentumly updates on pseudo-labels. Secondly, to mitigate catastrophic forgetting, we construct an episodic memory by distance-based sampling for representative samples and region-aware sampling for diverse samples, which are replayed by a knowledge distillation loss. These two modules mutually benefit each other to achieve a balanced perception on all classes—the disambiguation module ensures accurate training on new classes while the memory replay module strengthens the old classes. We combine PLL methods with various IL techniques to establish optimal baselines, demonstrating that PGDR achieves state-of-the-art results. In IPLL, our method outperforms the best baseline by 6.05% and 11.60% on the CIFAR100 and Tiny-ImageNet.
@@
The data appear in a sequence and the model continuously learns novel knowledge while maintaining the discrimination ability for previous knowledge. There is a classical incremental learning variant—blurry incremental learning (Bang et al. 2021; Moon et al. 2023), where different stages exhibit distinct data distributions. Due to the intersection of the label space, it faces heightened ambiguity. Nonetheless, IL and its variants are confronted with the challenge of knowledge forgetting (Kim, Park, and Han 2024; Li et al. 2024). To address this issue, existing methods can be categorized into three major strategies. For architecture-based methods (Yoon et al. 2020; Marouf et al. 2024), different model parameters are allocated for each task. The regularization-based methods (Szatkowski et al. 2023; Bian et al. 2024) introduce regularization terms into the objective function. The replay-based methods (Wang et al. 2022a; Bhat, Zonooz, and Arani 2023; Yoo et al. 2024) retain a subset of historical samples and incorporate them into subsequent tasks.
The IL variants discussed above all assume accurately labeled data. More and more researchers are increasingly interested in incremental learning in limited or unsupervised settings (Kang et al. 2022; Boschini et al. 2022; Wu et al. 2023), such as unsupervised continual learning (Cha, Cho, and Moon 2024), semi-supervised continual learning (Fan et al. 2024), and noisy labeled continual learning (Bang et al. 2022). However, current work overlooks label ambiguity, prompting us to explore incremental learning under ambiguous labels.
@@
The fundamental challenge in PLL is label disambiguation, requiring the model to select the true label from the set of candidate labels (Li, Li, and Ouyang 2020; Yan and Guo 2021; Wang, Zhang, and Li 2022; Bao, Rui, and Zhang 2024). Some PLL methods enhance candidate label disambiguation through adversarial learning (Zhang et al. 2020, 2023), and the majority of work (Wang et al. 2022c; Tian et al. 2024) is geared towards devising appropriate objectives for PLL. The strategy (Lv et al. 2023) based on averaging assigns equal weights to the labels in the candidate labels, and then obtains predictions by averaging the output. The disambiguation strategy based on identification (Yu and Zhang 2017; Tian et al. 2024) treats the true labels of samples as latent variables and iteratively optimizes the objective function of these latent variables. Based on self-training disambiguation strategies (He et al. 2022; Yan and Guo 2023; Dong et al. 2023; Liu et al. 2024), pseudo-labels are used as the model’s supervisory information for training, e.g., PRODEN (Lv et al. 2020) re-normalizes the classifier’s output and PiCO (Wang et al. 2022c) introduces contrastive learning. Despite the promise, existing PLL work mostly assumes a static data distribution, which is typically not available in practice. Therefore, we introduce incremental learning to investigate IPLL that is more suitable for real-world scenarios.
@@
The goal of IPLL is to sequentially learn a unified model from ambiguous supervised datasets and classify unseen test samples of all classes that have been learned so far. Formally, assume we are given a stream of datasets $\{{\cal D}_{t}\}_{t=1}^{T}$, where each subset ${\cal D}_{t}=\{{({\bm{x}_{i}^{t}},{{\cal S}_{i}^{t}})}\}_{i=1}^{{N_{t}}}$ contains $N_{t}$ samples. Here, ${\bm{x}^{t}_{i}}\in{\cal X}$ represents a sample in the input space $\mathbb{R}{{}^{d}}$. Different from the supervised setup where the ground truth $y_{i}$ is known, we follow the setup of previous PLL studies (Feng et al. 2020; Lv et al. 2020; Wang et al. 2022c) and allow the annotator to assign a rough candidate label set $\mathcal{S}_{i}^{t}\subset\mathcal{Y}_{t}$ containing the true label, i.e., $y_{i}^{t}\in\mathcal{S}_{i}^{t}$. For the label space, we consider a blurry incremental learning (Bang et al. 2021) that can be ubiquitous in real-world applications [^1] The classic IL setup simply assumes the label spaces have no intersection, which is less practical and the label ambiguity issue is typically not significant. , whose data distribution demonstrates: (i)-a majority of samples with new labels emerge incrementally $\mathcal{Y}_{t}=\mathcal{Y}_{t-1}\cup\mathcal{Y}^{\text{new}}_{t}$; (ii)-each subset ${{\cal D}_{t}}$ potentially contains data samples from all labels in $\mathcal{Y}_{t}$. At step $t$, the goal of IPLL is to train a model $f$ from the new dataset $\mathcal{D}_{t}$ without interfering with previous data, where $f$ consists of the feature extractor backbone $\phi$ with a fully-connected layer upon it. During training, since the ground-truth label is not accessible, we assign each sample $\bm{x}_{i}$ a pseudo-label vector $\bm{p}_{i}\in[0,1]^{|\mathcal{Y}_{t}|}$ and update the model by cross-entropy loss:
${\mathcal{L}_{\text{ce}}}=\frac{1}{{|\mathcal{Y}_{t}|}}\sum\limits_{j\in{\mathcal{Y}}_{t}}{-{p_{ij}}\log(f_{j}({\bm{x}_{i}}))}.$ (1)
@@
Next, we re-allocate the candidate label sets according to our data separation procedure,
@@
where $\bm{z}_{i}$ is a one-hot vector, i.e., ${z_{ij}}=\mathbb{I}(j=\arg\mathop{\max}_{j\in{{\cal S}_{i}^{t}}}{f_{j}}({\bm{x}_{i}}))$. Lastly, we show how our momentum updating procedure helps balance the class perception. At the beginning of task $t$, the relatively accurate old class can dominate the disambiguation procedure. Hence, we assign a large $\beta$ such that the supervision signals retain more probability on new classes. This helps to adversarially counteract the model’s bias towards the old classes through a procedure analogous to label smoothing (Müller, Kornblith, and Hinton 2019). As the prediction on new classes becomes more accurate (i.e., more balanced perception), we set a smaller $\beta$ value to facilitate the convergence of the pseudo-labels.
@@
Apart from those representative samples, we further collect a few samples that can better preserve the diversity of old class distribution. In our IPLL setup, this should be done more carefully due to the label ambiguity. We propose selecting a sample with two characteristics. First, it should lie in a smooth local manifold to avoid including falsely disambiguated samples. Second, it does not lie in a local region near those already selected samples and carries mutually different patterns. Formally speaking, we calculate the sum of distance measures from a sample to its $K$-nearest neighbor:
@@
In our implementation, for each class, we first select up to ${N_{d}}$ diverse samples and then, select $N_{r}=m/|{{\cal Y}_{t}}|-{N_{d}}$ representative samples, resulting in ${\cal M}^{t}={\cal M}^{t}_{r}\cup{\cal M}^{t}_{k}$. This provides richer class information for subsequent learning, mitigating the forgetting of old classes.
@@
${\mathcal{L}_{\text{kd}}}=\frac{1}{|\mathcal{Y}_{t-1}|}{\sum_{j\in\mathcal{Y}_{t-1}}-f^{old}_{j}({\bm{x}_{i}})\log(f_{j}({\bm{x}_{i}}))}.$ (10)
@@
Datasets. We perform experiments on CIFAR100 (Krizhevsky, Hinton et al. 2009) and Tiny-ImageNet (Le and Yang 2015). Additionally, we further conduct experiments on CUB200 (Welinder et al. 2010). For the experimental setup of IPLL, we reference the settings of PLL (Lv et al. 2020) and IL (Li and Hoiem 2016; Zhao et al. 2020). We generate partially labeled datasets by manually flipping negative labels $\bar{y}\neq y$ to false-positive labels with probability $q=P(\bar{y}\in{\cal Y}_{t}|\bar{y}\neq y)$. In the $t$-th task, all $|{\cal Y}_{t}|-1$ negative labels have a uniform probability to be false positive and we aggregate the flipped ones with the ground-truth to form the candidate set ${\cal S}_{i}^{t}$. The flipping probability $q$ is set at 0.1 and 0.2. We randomly partition all classes into 10 tasks, i.e., $T=10$. In the first task, there are exclusively new classes, while in the remaining $T-1$ tasks, a substantial number of samples belong to new classes, with relatively fewer samples belonging to old classes. To be specific, for a certain category of samples, $W\%$ of the samples emerge as new class samples in the $t$-th task, while the remaining $(1-W)\%$ of the samples uniformly appear as old class samples in each subsequent task $\{t+1,t+2,...,T\}$. We consider the degree of blending new and old class data to be “(1-$W$)-blurry” and set the value of $W$ to 90 (10-blurry) and 70 (30-blurry).
Baselines. We compare PGDR with two PLL methods: PiCO (Wang et al. 2022c) and PaPi (Xia et al. 2023). We also discuss PRODEN (Lv et al. 2020) in the appendix. We compare PGDR with five incremental learning methods: 1) iCaRL (Rebuffi et al. 2017); 2) BiC (Wu et al. 2019); 3) WA (Zhao et al. 2020); 4) ER-ACE (Caccia et al. 2022); 5) ANCL (Kim et al. 2023). To ensure experimental fairness and credibility, we combine these five incremental learning methods with PiCO (Wang et al. 2022c) and PaPi (Xia et al. 2023) to achieve basic label disambiguation. To objectively evaluate our method, we respectively plug the contrastive learning module of PiCO and the Kullback-Leibler divergence of PaPi into our method.
@@
In the experiment, we employ ResNet-18 for feature extraction. The network model is trained using SGD with a momentum of 0.9. Trained for 200 epochs on the CIFAR100, and trained for 300 epochs on the Tiny-ImageNet. The learning rate starts at 0.1 for PiCO and 0.01 for PaPi. Batch sizes are set to 256 and 128 for the CIFAR100 and Tiny-ImageNet, respectively, with a maximum sample storage limit of $m$ of 2000. For prototypes, a moving average coefficient $\gamma$ of 0.5 is used. In the sample selection for the storage phase, we set the number of nearest neighbors $K$ to 10, with a maximum limit $N_{d}$ for diverse sample storage of $0.67*m/|{\cal Y}_{t}|$. In the label disambiguation stage, the threshold ${\alpha}$ is set to 0.8. We linearly ramp down $\beta$ from 0.8 to 0.6 to ensure the full utilization of differential labels in the early stages of training. This helps mitigate the introduction of noisy information resulting from inaccurate early predictions. The experiments regarding PiCO and PaPi reference their experimental parameters.
@@
Comparison on fine-grained datasets.
When semantically similar classes are concentrated in a particular stage, disambiguation becomes more challenging. To address this, we reference (Wang et al. 2022c, b) and evaluate our method on two fine-grained datasets: 1) CIFAR100 dataset with hierarchical labels, i.e., CIFAR100-H. There are 5 categories under each superclass, and there is no intersection between different superclasses; 2) CUB200 (Welinder et al. 2010) dataset with 200 species of birds. For CIFAR100-H, the new classes in each task are derived from two unseen superclasses in IPLL. The flipping probabilities $q$ are 0.1 and 0.2. As shown in Table 2, all solutions exhibited a noticeable decrease in performance in the more challenging setting. On the CUB200 dataset, when $q$ is set to 0.1, the gaps between our method and the best baseline that incorporates PiCO are 2.67% (10-blurry) and 6.24% (30-blurry). This thoroughly validates the efficacy of our method, even under highly challenging data settings.
@@
Ablation Study
The additional introduction of the loss function in PGDR hinders visualizing the effects of each module, reducing the interpretability of ablation results. Thus, we present ablation results on PGDR without contrastive learning and Kullback-Leibler divergence to demonstrate our method’s effectiveness. More ablation experiments can be found in Appendix.
@@
To validate the effectiveness of the label disambiguation module, we choose the variants 1) PGDR w MP and 2) PGDR w PP; see Appendix for more details. The first variant utilizes the label computation method of PRODEN. The second variant utilizes PiCO’s labeling method. The Table 3 shows that our disambiguation solution is significantly better than the two variants. To explore further, we visualized category changes under the Tiny-ImageNet 10-blurry (Figure 6). PGDR w PP frequent updates to prototypes accumulate model bias, leading to a noticeable decline in classifications when new class data is introduced. In comparison, our solution alleviates knowledge confusion, balancing the model’s perception of categories, especially in old classes.

### v4

@@ Introduction
To reduce annotation costs, we study a novel weakly supervised learning framework dubbed incremental partial label learning (IPLL), where (i)-data is given sequentially as a stream; (ii)-each task contains a vast number of new class samples while potentially carries old class data, and (iii)-each sample is associated with a candidate label set instead of the ground truth; see Figure 1. Notably, since the annotator may confuse previous experience with the true label of the sample, our IPLL setup allows the candidate label set of the sample to include both new and old classes, while ensuring the inclusion of the true one. Arguably, the IPLL problem is deemed more practical in real-world scenarios due to its relatively lower cost to annotations.
The key to successful learning from ambiguous supervision is label disambiguation, i.e., identifying the true labels from the candidate sets. To achieve this, existing partial label learning (PLL) algorithms (Wang et al. 2022b; Yan and Guo 2023; Li et al. 2023; Jia et al. 2024; Liu et al. 2024) mostly rely on the self-training algorithm that assigns pseudo-labels using the model predictions. However, such a strategy can be problematic in our IPLL framework. On the one hand, the dynamically changing environment leads to the forgetting of past knowledge, which amplifies knowledge confusion that hinders the disambiguation procedure. On the other hand, ambiguous supervision significantly undermines the ability of DNNs to acquire meaningful knowledge from samples, exacerbating catastrophic forgetting. As exemplified in our empirical studies, directly combining the PLL method with the IL method also leads to inferior results. Specifically, when combining the most popular PLL method PiCO (Wang et al. 2022b) with one classic IL method iCaRL (Rebuffi et al. 2017), the model overly emphasizes old classes in the early stages of each task (see Figure 2(a)), thus exacerbating the confusion of old class knowledge (see Figure 2(b)). In the more challenging Tiny-ImageNet, the negative impacts of classification bias become more pronounced (see Figure 10(c)). This gives rise to the fundamental issue in IPLL—how to balance the model’s perception of new and old knowledge.
@@
Figure 2: (a) Class distribution based on both predicted labels and ground-truth labels. The real/estimated distribution of old classes (0$\sim$9) and new classes (10$\sim$19) in the early stage of the second task of the CIFAR100 (10 tasks). Combining the PLL baseline (PiCO) and the IL baseline (iCaRL) leads to classification bias. (b) Comparison of our PGDR and PiCO with iCaRL in IPLL. It displays the average accuracy of new and old classes for each task in CIFAR100 where PiCO with iCaRL demonstrates inferior and unstable performance compared to our method.
To address this issue, we propose the Prototype-Guided Disambiguation and Replay Algorithm (PGDR), which leverages class prototypes carrying rich category information as proxies to balance the model’s perception of classes. Concretely, to alleviate label confusion, we propose a prototype-guided label disambiguation strategy, which firstly performs distance-based old/new sample separation and then, momentumly updates on pseudo-labels. Secondly, to mitigate catastrophic forgetting, we construct an episodic memory by distance-based sampling for representative samples and region-aware sampling for diverse samples, which are replayed by a knowledge distillation loss. These two modules mutually benefit each other to achieve a balanced perception on all classes—the disambiguation module ensures accurate training on new classes while the memory replay module strengthens the old classes. We conduct comprehensive experiments on benchmark datasets to show that PGDR establishes state-of-the-art performance. In IPLL, our method outperforms the best baseline by 6.05% and 11.60% on the CIFAR100 and Tiny-ImageNet.
@@
The data appear in a sequence and the model continuously learns novel knowledge while maintaining the discrimination ability for previous knowledge. There is a classical incremental learning variant—blurry incremental learning (Bang et al. 2021; Moon et al. 2023), where different stages exhibit distinct data distributions. Due to the intersection of the label space, it faces heightened ambiguity. Nonetheless, IL and its variants are confronted with the challenge of knowledge forgetting (Kim, Park, and Han 2024; Li et al. 2024). To address this issue, existing methods can be categorized into three major strategies. For architecture-based methods (Yoon et al. 2020; Marouf et al. 2024), different model parameters are allocated for each task. The regularization-based methods (Bian et al. 2024) introduce regularization terms into the objective function. The replay-based methods (Bhat, Zonooz, and Arani 2023; Yoo et al. 2024) retain a subset of historical samples and incorporate them into subsequent tasks. Moreover, while some works (Lange and Tuytelaars 2021; Asadi et al. 2023) employ prototypes, they mainly focus on using them to alleviate knowledge forgetting and do not investigate their effectiveness in other aspects, distinguishing our approach from theirs.
The IL variants discussed above all assume accurately labeled data. More and more researchers are increasingly interested in incremental learning in limited or unsupervised settings (Wu et al. 2023), such as unsupervised continual learning (Cha, Cho, and Moon 2024), semi-supervised continual learning (Fan et al. 2024), and noisy labeled continual learning (Bang et al. 2022). However, most of the work overlooks label ambiguity. Although (Yu, Wang, and Zhang 2024) attempts to address this issue, the subsequent samples are unlabeled, making it similar to a semi-supervised learning problem.
@@
The fundamental challenge in PLL is label disambiguation, requiring the model to select the true label from the set of candidate labels (Yan and Guo 2021; Lyu et al. 2021; Wang, Zhang, and Li 2022; Bao, Rui, and Zhang 2024). Some PLL methods enhance candidate label disambiguation through adversarial learning (Zhang et al. 2020, 2023), and the majority of work (Wang et al. 2022b; Tian et al. 2024) is geared towards devising appropriate objectives for PLL. The strategy (Lv et al. 2023) based on averaging assigns equal weights to the labels in the candidate labels, and then obtains predictions by averaging the output. The disambiguation strategy based on identification (Yu and Zhang 2017; Tian et al. 2024) treats the true labels of samples as latent variables and iteratively optimizes the objective function of these latent variables. Based on self-training disambiguation strategies (He et al. 2022; Yan and Guo 2023; Dong et al. 2023; Liu et al. 2024), pseudo-labels are used as the model’s supervisory information for training, e.g., PRODEN (Lv et al. 2020) re-normalizes the classifier’s output and PiCO (Wang et al. 2022b) introduces contrastive learning. Despite the promise, existing PLL work mostly assumes a static data distribution, which is typically not available in practice. Therefore, we introduce incremental learning to investigate IPLL that is more suitable for real-world scenarios.
@@
The goal of IPLL is to sequentially learn a unified model from ambiguous supervised datasets and classify unseen test samples of all classes that have been learned so far. Formally, assume we are given a stream of datasets $\{{\cal D}_{t}\}_{t=1}^{T}$, where each subset ${\cal D}_{t}=\{{({\bm{x}_{i}^{t}},{{\cal S}_{i}^{t}})}\}_{i=1}^{{N_{t}}}$ contains $N_{t}$ samples. Here, ${\bm{x}^{t}_{i}}\in{\cal X}$ represents a sample in the input space $\mathbb{R}{{}^{d}}$. Different from the supervised setup where the ground truth $y_{i}$ is known, we follow the setup of previous PLL studies (Feng et al. 2020; Lv et al. 2020; Wang et al. 2022b) and allow the annotator to assign a rough candidate label set $\mathcal{S}_{i}^{t}\subset\mathcal{Y}_{t}$ containing the true label, i.e., $y_{i}^{t}\in\mathcal{S}_{i}^{t}$. For the label space, we consider a blurry incremental learning (Bang et al. 2021) that can be ubiquitous in real-world applications [^1] The classic IL setup simply assumes the label spaces have no intersection, which is less practical and the label ambiguity issue is typically not significant. , whose data distribution demonstrates: (i)-a majority of samples with new labels emerge incrementally $\mathcal{Y}_{t}=\mathcal{Y}_{t-1}\cup\mathcal{Y}^{\text{new}}_{t}$; (ii)-each subset ${{\cal D}_{t}}$ potentially contains data samples from all labels in $\mathcal{Y}_{t}$. At step $t$, the goal of IPLL is to train a model $f$ from the new dataset $\mathcal{D}_{t}$ without interfering with previous data, where $f$ consists of the feature extractor backbone $\phi$ with a fully-connected layer upon it. During training, since the ground-truth label is not accessible, we assign each sample $\bm{x}_{i}$ a pseudo-label vector $\bm{p}_{i}$ and update the model by cross-entropy loss:
${\mathcal{L}_{\text{ce}}=-\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\sum_{j=1}^{{|\mathcal{Y}}_{t}|}p_{ij}\log(f_{j}(\bm{x}_{i}))}$ (1)
@@
We re-allocate the candidate label sets according to our data separation procedure,
@@
where $\bm{z}_{i}$ is a one-hot vector, i.e., ${z_{ij}}=\mathbb{I}(j=\arg\mathop{\max}_{j\in{{\cal S}_{i}^{t}}}{f_{j}}({\bm{x}_{i}}))$, and $\beta$ is a hyperparameter. Practically, we assign a large $\beta$ at the beginning of each task due to the unreliable model prediction of new classes. As the prediction on new classes becomes more accurate, we set a smaller $\beta$ value to facilitate the convergence of the pseudo-labels.
@@
Apart from those representative samples, we further collect a few samples that can better preserve the diversity of old class distribution. In our IPLL setup, this should be done more carefully due to the label ambiguity. We propose selecting a sample with two characteristics. First, it should lie in a smooth local manifold to avoid including falsely disambiguated samples. Second, it does not lie in a local region near those already selected samples and carries mutually different patterns. Formally speaking, we calculate the sum of distance measure from a sample to its $K$-nearest neighbor:
@@
In our implementation, for each class, we first select up to ${N_{d}}$ diverse samples and then, select $N_{r}=m/|{{\cal Y}_{t}}|-{N_{d}}$ representative samples, resulting in ${\cal M}^{t}={\cal M}^{t}_{r}\cup{\cal M}^{t}_{k}$. This provides richer class information for subsequent learning, mitigating the forgetting of old classes. ${\mathcal{L}_{\text{ce}}}$ also includes this portion of samples
@@
${\mathcal{L}_{\text{kd}}}=-\frac{1}{|\mathcal{D}^{\prime}_{t}|}\sum_{i=1}^{|\mathcal{D}^{\prime}_{t}|}\sum_{j=1}^{{|\mathcal{Y}}_{t-1}|}{f^{old}_{j}({\bm{x}_{i}})\log(f_{j}({\bm{x}_{i}}))}.$ (10)
@@
Datasets.
We perform experiments on CIFAR100 (Krizhevsky, Hinton et al. 2009) and Tiny-ImageNet (Le and Yang 2015). Additionally, we further conduct experiments on CUB200 (Welinder et al. 2010). For the experimental setup of IPLL, we reference the settings of PLL (Lv et al. 2020) and IL (Li and Hoiem 2016; Zhao et al. 2020). We generate partially labeled datasets by manually flipping negative labels $\bar{y}\neq y$ to false-positive labels with probability $q=P(\bar{y}\in{\cal Y}_{t}|\bar{y}\neq y)$. In the $t$-th task, all $|{\cal Y}_{t}|-1$ negative labels have a uniform probability to be false positive and we aggregate the flipped ones with the ground-truth to form the candidate set ${\cal S}_{i}^{t}$. The flipping probability $q$ is set at 0.1 and 0.2. We randomly partition all classes into 10 tasks, i.e., $T=10$. In the first task, there are exclusively new classes, while in the remaining $T-1$ tasks, a substantial number of samples belong to new classes, with relatively fewer samples belonging to old classes. To be specific, for a certain category of samples, $W\%$ of the samples emerge as new class samples in the $t$-th task, while the remaining $(100-W)\%$ of the samples uniformly appear as old class samples in each subsequent task $\{t+1,t+2,...,T\}$. We consider the degree of blending new and old class data to be “(100-$W$)-blurry” and set the value of $W$ to 90 (10-blurry) and 70 (30-blurry).
Baselines.
We compare PGDR with two PLL methods: PiCO (Wang et al. 2022b) and PaPi (Xia et al. 2023). We also discuss PRODEN (Lv et al. 2020) in Appendix. We compare PGDR with five incremental learning methods: 1) iCaRL (Rebuffi et al. 2017); 2) BiC (Wu et al. 2019); 3) WA (Zhao et al. 2020); 4) ER-ACE (Caccia et al. 2022); 5) ANCL (Kim et al. 2023). To ensure experimental fairness and credibility, we combine these five incremental learning methods with PiCO and PaPi to achieve basic label disambiguation. To objectively evaluate our method, we respectively plug the contrastive learning module of PiCO and the Kullback-Leibler divergence of PaPi into our method.
@@
We employ ResNet-18 for feature extraction. The network model is trained using SGD with a momentum of 0.9. Trained for 200 epochs on the CIFAR100, and trained for 300 epochs on the Tiny-ImageNet. The learning rate starts at 0.1 for PiCO and 0.01 for PaPi. Batch sizes are set to 256 and 128 for the CIFAR100 and Tiny-ImageNet, with a maximum sample storage limit of $m$ of 2000. For prototypes, a moving average coefficient $\gamma$ of 0.5 is used. In the sample selection for the storage phase, we set the number of nearest neighbors $K$ to 10, with a maximum limit $N_{d}$ for diverse sample storage of $0.67*m/|{\cal Y}_{t}|$. In the label disambiguation stage, the threshold ${\alpha}$ is set to 0.8. We linearly ramp down $\beta$ from 0.8 to 0.6 to ensure the full utilization of differential labels in the early stages of training. This helps mitigate the introduction of noisy information resulting from inaccurate early predictions. The experiments regarding PiCO and PaPi reference their experimental parameters.
@@
Comparison on fine-grained datasets. When semantically similar classes are concentrated in a particular stage, disambiguation becomes more challenging. To address this, we reference (Wang et al. 2022b, a) and evaluate our method on two fine-grained datasets: 1) CIFAR100 dataset with hierarchical labels, i.e., CIFAR100-H. There are five categories under each superclass, and there is no intersection between different superclasses; 2) CUB200 (Welinder et al. 2010) dataset with 200 species of birds. For CIFAR100-H, the new classes in each task are derived from two unseen superclasses in IPLL. The flipping probabilities $q$ are 0.1 and 0.2. As shown in Table 2, all solutions exhibited a noticeable decrease in performance in the more challenging setting. On the CUB200 dataset, when $q$ is set to 0.1, the gaps between our method and the best baseline that incorporates PiCO are 2.67% (10-blurry) and 6.24% (30-blurry). This thoroughly validates the efficacy of our method, even under highly challenging data settings.
Ablation Study
The additional introduction of the loss function in PGDR hinders visualizing the effects of each module, reducing the interpretability of ablation results. Thus, we present ablation results on PGDR without contrastive learning and Kullback-Leibler divergence to demonstrate our method’s effectiveness. More experiments can be found in Appendix.
@@
To validate the effectiveness of the label disambiguation module, we choose the variants 1) PGDR w MP and 2) PGDR w PP; see Appendix. The first variant utilizes the PRODEN label computation method. The second variant utilizes PiCO’s labeling method. The Table 3 shows that our disambiguation solution is significantly better than the two variants. To explore further, we visualized category changes under the Tiny-ImageNet 10-blurry (Figure 10). PGDR w PP frequent updates to prototypes accumulate model bias, leading to a noticeable decline in classifications when new class data is introduced. In comparison, our solution alleviates knowledge confusion, balancing the model’s perception of categories, especially in old classes.